# EvalAlign: Supervised Fine-Tuning Multimodal LLMs with Human-Aligned Data for Evaluating Text-to-Image Models

### v2

@@ Abstract
The recent advancements in text-to-image generative models have been remarkable. Yet, the field suffers from a lack of evaluation metrics that accurately reflect the performance of these models, particularly lacking fine-grained metrics that can guide the optimization of the models. In this paper, we propose EvalAlign, a metric characterized by its accuracy, stability, and fine granularity. Our approach leverages the capabilities of Multimodal Large Language Models (MLLMs) pre-trained on extensive datasets. We develop evaluation protocols that focus on two key dimensions: image faithfulness and text-image alignment. Each protocol comprises a set of detailed, fine-grained instructions linked to specific scoring options, enabling precise manual scoring of the generated images. We Supervised Fine-Tune (SFT) the MLLM to align closely with human evaluative judgments, resulting in a robust evaluation model. Our comprehensive tests across 24 text-to-image generation models demonstrate that EvalAlign not only provides superior metric stability but also aligns more closely with human preferences than existing metrics, confirming its effectiveness and utility in model assessment.
@@ 1 Introduction
Text-to-image models, such as DALL·E series [38, 37, 3], Imagen [43], and Stable Diffusion [33], have significantly impacted various domains such as entertainment, design, and education, by enabling high-quality image generation. These technologies not only advance the field of text-to-image generation but also enhance related applications such as video generation [4, 60], image editing [46, 18, 59], and personalized image generation [13, 42]. Despite this prevalence, in the practical use of generative models, we still lack evaluation methods to accurately evaluate existing text-to-image models.
Existing benchmarks for text-to-image generation models are neither comprehensive nor accurate enough to meet current evaluation needs. The reasons for this include: (1) Limited model parameters: Current evaluation models have too few parameters, which restricts their ability to accurately represent images, leading to significant discrepancies compared to human evaluations. (2) Training data limitations: Some evaluation methods, such as Inception Score (IS) [44], Frechet Inception Distance (FID) [15], and CLIP Score [14], use models that have not been trained with synthesized images, which may introduce training bias and flaws the evaluation. (3) High annotation costs: Some methods, such as ImageReward [56], HPS [55] and HPS v2 [54], rely heavily on extensive human annotations, which significantly increases the cost of labeling. (4) Lack of detailed evaluation metric: The evaluation metrics do not provide fine-grained interpretability, preventing them from guiding model optimization effectively. (5) Computational inefficiency: The evaluation models require substantial computational resources, making them inefficient.
To address these issues, we propose EvalAlign, which offers low-cost, accurate, and efficient model evaluations while providing fine-grained, interpretable metrics. Specifically, we Supervised Fine-Tune (SFT) a Multimodal Large Language Model (MLLM) to align it with human annotations. Our focus is on two key evaluation aspects: image faithfulness and text-image alignment.
Due to its pre-training on large-scale datasets and a large number of model parameters, the MLLM demonstrates excellent understanding and generalization capabilities for images and instructions. This makes it possible to design evaluation instruction based on MLLM to evaluate text-to-image models. However, since the pre-trained datasets do not include model-generated images (such as distorted body structures and human hands) or evaluation-related text instructions, using MLLM directly for model evaluation does not yield optimal results. Therefore, we employ SFT on a small amount of manually annotated data to align MLLM with human annotations for text-to-image generation. In summary, our main contributions can be summarized as follows:
We introduce a dataset specifically designed to address the evaluation challenges of text-to-image models. This dataset, derived from multiple data sources, has been thoroughly cleaned and systematically annotated by human. It enables precise evaluation of text-image alignment and image faithfulness.
@@
We propose the EvalAlign evaluation metric, which accurately aligns evaluation metrics with human preferences. This method is cost-effective in terms of annotation and training, computationally efficient, and provides interpretable metrics.
@@
We conduct evaluations over 24 text-to-image models and compare EvalAlign with existing evaluation methods. Extensive experiments demonstrate that EvalAlign outperforms other methods in evaluating model performance.
@@ 2 Related Work
Evaluations and benchmarks of text-to-image generation. Despite the incredible progress achieved by text-to-image generation, evaluations and benchmarks in this area are far from flawless and contain critical limitations. For example, the most commonly used metrics, IS [44], FID [15], and CLIPScore [14] are broadly recognized as inaccurate for their inconsistency with human perception. To address, HPS series [55, 54], PickScore [20], and ImageReward [56] introduced human preference prior on image assessing to the benchmark, thereby allowing better correlation with image quality. However, with varying source and size of training data, these methods merely score the evaluated images in a coarse and general way, which cannot serve as an indication for model evolution. Meanwhile, HEIM [21] combined automatic and human evaluation and holistically evaluated text-to-image generation in 12 aspects, such as alignment, toxicity, and so on. As a consequence, HEIM relies heavily on human labour, limiting its application within budget-limited research groups severely. [30] standardized the protocol and settings of human evaluation, ensuring its verifiable and reproducible. There are also some works bear a resemble with us. For instance, TIFA [16], Gecko [53] and LLMScore [27] also formulate the evaluation as a set of visual question answering procedure and use LLMs as evaluation models. However, while they all mainly focus on text-image alignment, our approach takes both text-image alignment and image faithfulness into consideration. Moreover, the evaluation of LLMScore requires an object detection stage, which introduces significantly extra inference latency to the evaluation pipeline.
@@
Inception Score [44] | NeurIPS 2016 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
FID [15] | NeurIPS 2017 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
CLIP-score [14] | EMNLP 2021 | ✗ | ✗ | ✓ | 400M | 400M | – | ✗ | ✓ |
HPS [55] | ICCV 2023 | ✓ | ✗ | ✓ | 25K | 98K | 25K | – | – |
TIFA [16] | ICCV 2023 | ✓ | ✓ | ✓ | 4K | – | 25K | ✗ | ✓ |
TVRHE [30] | CVPR 2023 | ✓ | ✗ | ✗ | – | – | – | ✓ | ✗ |
ImageReward [56] | NeurIPS 2023 | ✓ | ✗ | ✓ | 8.8K | 68K | 137K | – | – |
PickScore [20] | NeurIPS 2023 | ✓ | ✗ | ✓ | 35K | 1M | 500K | – | – |
HPS v2 [54] | arXiv 2023 | ✓ | ✗ | ✓ | 107K | 430K | 645K | – | – |
HEIM [21] | NeurIPS 2023 | ✓ | ✓ | ✗ | – | – | – | ✓ | ✓ |
Gecko [53] | arXiv 2024 | ✓ | ✓ | ✓ | 2K | – | 108K | ✗ | ✓ |
LLMScore [27] | arXiv 2024 | ✓ | ✓ | ✓ | – | – | – | ✗ | ✓ |
@@
Multimodal Large Language Models (MLLMs). Pretrained on massive text-only and image-text data, MLLMs have exhibited exceptional image-text joint understanding and generalization abilities, facilitating a large spectrum of downstream applications. Among the works major in MLLMs, LLaVA [26, 24] and MiniGPT4 [61, 5] proposed to conduct visual instruction tuning during SFT, so that MLLMs can be easily aligned to human preference and precisely answer fine-grained questions on visual content. Meanwhile, Video-LLaMA [58] and VideoChat [22] utilized MLLMs for video understanding. VILA [23] quantitatively proved that involving text-only instruction-tuning data during SFT can further ameliorate model performance on text-only and multimodal downstream tasks. LLaVA-NeXT [25] extracted visual tokens for both the resized input image and the segmented sub-images to provide more detailed visual information for MLLMs, achieving significant performance bonus on tasks with high-resolution input images.
In this paper, we manually annotate a minimal amount of high-quality visual instruction-tuning data to adapt MLLMs to perform human preference aligned and fine-grained evaluation. Owing to the generalization and multimodal understanding abilities of MLLMs, our experiments demonstrated that the finetuned MLLMs can accurately assess text-to-image generation models by following instructions regarding various fine-grained criteria of synthesised image.
@@ 3 EvalAlign Dataset Construction
To train, validate and test the effectiveness of our evaluation models, we build EvalAlign benchmark. Specifically, EvalAlign dataset is a meticulously annotated collection featuring fine-grained annotations for images generated from text prompts. This dataset comprises 21k images, each accompanied by detailed instructions. The compilation process for the EvalAlign Dataset encompasses prompt collection, image generation, and precise instruction-based annotation.
@@ 3.1 Prompts and Images Collection
Prompt collection. To assess the capabilities of our model in terms of text-image alignment and image faithfulness, we collect, filter, and clean prompts from existing evaluation datasets and generated prompts based on LLM. These prompts encompass a diverse range from real-world user prompts, prompts generated through rule-based templates with LLM, to manually crafted prompts. Specifically, the utilized prompts are soureced from HPS [55], HRS-Bench [2], HPSv2 [54], TIFA [16], DSG [7], T2I-Comp [17], Winoground [50], DALL-EVAL [8], DiffusionDB [52], PartiPrompts [57], DrawBench [43], and JourneryDB [49].
Prompt curation. The prompts are collected for facilitating fine-grained and comprehensive evaluation on text-to-image models in terms of image faithfulness and text-image alignment. Considering some of the collected prompt are unsuitable for these two evaluation tasks, we filter the collected prompts to ensure their quantity, quality and diversity. For the image faithfulness evaluation task, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment. Our filtering process initially selects prompts that describe human, animals, and other real objects. After deduplicating these prompts, we carefully selected 1,500 distinct prompts with varying topic, background and style for further evaluation. The selected prompts encompass 10k subjects across 15 categories. For the text-image alignment task, we refined our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts. Only those prompts containing relevant descriptions and exceeding 15 words in length were considered, culminating in a final set of 1,500 prompts for evaluation.
Image generation. To train and evaluate the MLLM, we use a diverse set of images generated by various models using the aforementioned prompts, allowing for detailed human annotation. For each prompt, multiple images are generated across different models. The dataset is enriched by images synthesized by models with varying architectures and scales, enhancing the diversity essential for a comprehensive evaluation.
This variety not only tests the MLLM generalization capabilities but also aids in developing a model with broader applicability. The training dataset incorporates images from 8 models, whereas the test dataset spans 24 models. The inclusion of 16 unseen models in the test set is crucial for evaluating the MLLM to generalize beyond its training data. For detailed information on the inference setting of each model, please refer to the supplementary material. This structured approach ensures a robust framework for training and validating the MLLM, positioning it as a versatile and adaptive tool in the field of image generation from textual descriptions.
@@ 3.2 Data Annotation
Prompt annotation. For text prompts focused on text-image alignment, we begin by annotating the entities and their attributes within the text, as illustrated in Figure 1. Our annotators extract the entities mentioned in the prompts and label each entity with corresponding attributes, including quantity, color, spatial relationships, and actions. During the annotation, we also ask the annotators annotate the overall style of the image if described in the corresponding prompt and report prompts that contain toxic and NSFW content. These high-quality and detailed annotations facilitate the subsequent SFT training and evaluation of the MLLM. This meticulous annotation procedure ensures that the MLLM can accurately align and respond to the nuanced details specified in the prompts, enhancing both the training process and the model’s performance in generating images that faithfully reflect the described attributes and style.
@@
Image annotation. The images generated by text-to-image models often present challenges such as occluded human body parts, which can impede the effectiveness of SFT training and evaluation of the MLLM. Therefore, to address these challenges and enhance the model’s training and evaluative capabilities, specific annotations are applied to all images depicting human and animals. These annotations include: presence of human or animal faces; visibility of hands; visibility of limbs. By implementing these annotations, we ensure that the MLLM can more effectively learn from and assess the completeness and faithfulness of the generated images. This structured approach to annotation not only aids in identifying common generation errors but also optimizes the model’s ability to generate more accurate and realistic images, thereby improving both training outcomes and the model’s overall performance in generating coherent and contextually appropriate visual content.
@@
Instruction-Fine-tuning data annotation. To align the MLLM with human preference prior on fine-grained image assessing, we can train the model on a minimal amount of fine-grained human preference data through SFT training. As a consequence, we devise two sets of questions, each is concentrated on a specific fine-grained aspect of image faithfulness and image-text alignment, and ask human annotators to answer these questions to acquire the fine-grained human preference data. To aid the human annotators to understand the meaning of each question and its available answer options, thereby ensuring high annotation quality, we employ a thorough and comprehensive procedure of annotation preparation. First, we write a detailed annotation guideline and conduct a training for the annotators to explain the annotation guideline and answer their questions about the annotation (see supplementary materials for the annotation guideline). Then, we conduct a multi-turn trial annotation on another 50 synthesized images. After each trial, we calculate the Cohen’s kappa coefficient and interpret annotation guidelines for our annotators. In total, we conduct nine turns of trial annotation, and in the last turn of the trial, the Cohen’s kappa coefficient of our annotators reaches , indicating high inter-annotator reliability and high annotation quality. More results about the trial annotation will be included in the supplementary materials.
@@
After completing the aforementioned preparations, we delegate the images filtered during image annotation to 10 annotators and ask them to complete the annotation just as how they did in the trial annotation. Furthermore, during the whole annotation procedure, four text-to-image generation experts conduct random sampling quality inspection on the present annotated results, causing a second and a third re-annotation on 423 and 112 inspection-failed samples. Overall, owing to the valuable work of our human annotators and our fastidious annotation procedure, we get quality-sufficient instruction-tuning data required for the SFT training of the MLLM. More details of the annotation procedure will be introduced in the supplementary files.

### v3

@@ Abstract
The recent advancements in text-to-image generative models have been remarkable. Yet, the field suffers from a lack of evaluation metrics that accurately reflect the performance of these models, particularly lacking fine-grained metrics that can guide the optimization of the models. In this paper, we propose EvalAlign, a metric characterized by its accuracy, stability, and fine granularity. Our approach leverages the capabilities of Multimodal Large Language Models (MLLMs) pre-trained on extensive data. We develop evaluation protocols that focus on two key dimensions: image faithfulness and text-image alignment. Each protocol comprises a set of detailed, fine-grained instructions linked to specific scoring options, enabling precise manual scoring of the generated images. We supervised fine-tune (SFT) the MLLM to align with human evaluative judgments, resulting in a robust evaluation model. Our evaluation across 24 text-to-image generation models demonstrate that EvalAlign not only provides superior metric stability but also aligns more closely with human preferences than existing metrics, confirming its effectiveness and utility in model assessment.
@@ 1 Introduction
Text-to-image models, such as DALL·E series (Ramesh et al., 2022; Betker et al., 2023), Imagen (Saharia et al., 2022), and Stable Diffusion (Podell et al., 2023), have significantly impacted various domains such as entertainment, design, and education, by enabling high-quality image generation. These technologies not only advance the field of text-to-image generation but also bloom applications such as video generation (Blattmann et al., 2023; Zhang et al., 2023d; Tan et al., 2024b), image editing (Song et al., 2021; Huang et al., 2023b; Zhang et al., 2023c) and human image generation (Wang et al., 2024). Despite achieving incredible progress, the evaluation methods in this area are far from flawless and suffer heavily from data bias, as they are mainly trained on real images but are employed to evaluate synthesized images.
Since human-based evaluations are considerably costly in money and time, existing evaluation methods are primarily based on pretrained models, which are trained on real images. However, the trained real images are generated by humans and high in image faithfulness and text-image alignment because of their generation essence. Meanwhile, the evaluated images are synthesized by text-to-image models and encounter problems such as low image faithfulness or text-image alignment, constrained by the performance of generative models.
We dub the gap between the training data and the evaluated data as data bias, which may cause the evaluation models perform ill-suited on text-to-image evaluation. Because of the data bias, existing text-to-image evaluation methods performs poorly in synthesized image evaluations. Unfortunately, during our preliminary observation, nearly every synthesized images contain visual elements with low image faithfulness or text-image alignment, emphasize their significance on evaluation performance. Notably, there are also some works such as HPSv2 (Wu et al., 2023b) and PickScore Kirstain et al. (2024), where their evaluation models are trained synthesized images. However, in their evaluation settings, the utilized synthesized images are treated as real images as they don’t explicitly recognize the problem of synthesized images with low image faithfulness.
In view of these issues, we propose EvalAlign, a comprehensive, fine-grained and interpretable metric on text-to-image model assessing with low cost but high accuracy. To build EvalAlign, we first curate a dataset composed of fine-grained human feedback scores on synthesized images, with consideration of the corresponding prompts. The granularity of the feedback covers 11 skills categorized into two aspects: image faithfulness and text-image alignment. After that, we Supervised finetune (SFT) a Multimodal Large Language Model (MLLM) on the annotated dataset, aligning it with human prior on detailed and accurate text-to-image evaluation.
Owing to extensive pre-training and large model capacity, MLLMs demonstrate excellent image-text understanding and generalization capabilities. However, since the pre-training data does not include synthesized images with low image faithfulness or evaluation-related text instructions, using MLLMs directly for model evaluation may yield non-optimal results. Especially, we want to use MLLMs to support comprehensive and detailed evaluations, encompassing 11 skills and 2 aspects. The definitions and nuances of these may not be fully understood by the MLLM. Therefore, we employ SFT on a small amount of high-quality annotated data to align the MLLM with human judgement on evaluating synthesized images in criteria of 11 skills and 2 aspects. Notably, since the main intelligence of EvalAlign stems from the annotated dataset and the utilized MLLM, we will make them accessible to the public.
In summary, our main contributions can be summarized as follows:
We build a detailed human feedback dataset specifically designed to address the aforementioned challenges of text-to-image model evaluations. The annotated dataset is thoroughly cleaned, carefully balanced in topics, and systematically annotated by human. The dataset is composed by fine-grained human prior on evaluating synthesized images in criteria of 11 skills and 2 aspects.
@@
We propose EvalAlign, a text-to-image evaluation method which accurately aligns evaluation models with fine-grained human prior using the annotated dataset. EvalAlign exclusively supports an accurate, comprehensive, fine-grained and interpretable text-to-image evaluations. Besides EvalAlign is cost-effective in terms of annotation and training and computationally efficient.
@@
With EvalAlign, we conduct evaluations over 24 text-to-image models and compare EvalAlign with existing evaluation methods. Quantitative and qualitative experiments demonstrate that EvalAlign outperforms other methods in evaluating model performance.
@@ 2 Related Work
2.1 Benchmarks of text-to-image generation
Despite the incredible progress achieved by text-to-image generation Zhang et al. (2023a); Tan et al. (2024a), evaluations and benchmarks in this area are far from flawless and contain critical limitations. For example, the most commonly used metrics, IS (Salimans et al., 2016), FID (Heusel et al., 2017), and CLIPScore (Hessel et al., 2021) are broadly recognized as inaccurate for their inconsistency with human perception. To address, HPS series (Wu et al., 2023b; a), PickScore (Kirstain et al., 2024), and ImageReward (Xu et al., 2024) introduced human preference prior on image assessing to the benchmark, thereby allowing better correlation with image quality. However, with varying source and size of training data, these methods merely score the evaluated images in a coarse and general way, which cannot serve as an indication for model evolution. Meanwhile, HEIM (Lee et al., 2024) combined automatic and human evaluation and holistically evaluated text-to-image generation in 12 aspects, such as alignment, toxicity, and so on. As a consequence, HEIM relies heavily on human labour, limiting its application within budget-limited research groups severely. Otani et al. (2023) standardized the protocol and settings of human evaluation, ensuring its verifiable and reproducible. Considering the issues of existing benchmarks, we propose EvalAlign to offer a cost-efficient, comprehensive and fine-grained text-to-image model evaluation. Through our observations, we found that image faithfulness and text-image alignment are two key factors for comprehensive evaluation. Image faithfulness requires the model to generate visual elements that are consistently faithful to the real-world. For example, visual elements such as distorted body. Meanwhile, text-image alignment measures how the generated images are aligned with their corresponding prompts.
There are also some works bear a resemble with us. For instance, TIFA (Hu et al., 2023), Gecko (Wiles et al., 2024) and LLMScore (Lu et al., 2024) also formulate the evaluation as a set of visual question answering procedure and use LLMs as evaluation models. However, while they all mainly focus on text-image alignment, our approach takes both text-image alignment and image faithfulness into consideration. Moreover, the evaluation of LLMScore requires an object detection stage, which introduces significantly extra inference latency to the evaluation pipeline.
@@ 2.1 Benchmarks of text-to-image generation
Inception Score (Salimans et al., 2016) | NeurIPS 2016 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
FID (Heusel et al., 2017) | NeurIPS 2017 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
CLIP-score (Hessel et al., 2021) | EMNLP 2021 | ✗ | ✗ | ✓ | 400M | 400M | – | ✗ | ✓ |
HPS (Wu et al., 2023b) | ICCV 2023 | ✓ | ✗ | ✓ | 25K | 98K | 25K | – | – |
TIFA (Hu et al., 2023) | ICCV 2023 | ✓ | ✓ | ✓ | 4K | – | 25K | ✗ | ✓ |
TVRHE (Otani et al., 2023) | CVPR 2023 | ✓ | ✗ | ✗ | – | – | – | ✓ | ✗ |
ImageReward (Xu et al., 2024) | NeurIPS 2023 | ✓ | ✗ | ✓ | 8.8K | 68K | 137K | – | – |
PickScore (Kirstain et al., 2024) | NeurIPS 2023 | ✓ | ✗ | ✓ | 35K | 1M | 500K | – | – |
HPS v2 (Wu et al., 2023a) | arXiv 2023 | ✓ | ✗ | ✓ | 107K | 430K | 645K | – | – |
HEIM (Lee et al., 2024) | NeurIPS 2023 | ✓ | ✓ | ✗ | – | – | – | ✓ | ✓ |
Gecko (Wiles et al., 2024) | arXiv 2024 | ✓ | ✓ | ✓ | 2K | – | 108K | ✗ | ✓ |
LLMScore (Lu et al., 2024) | arXiv 2024 | ✓ | ✓ | ✓ | – | – | – | ✗ | ✓ |
@@
2.2 Multimodal Large Language Models (MLLMs)
Pre-trained on massive text-only and image-text data, MLLMs have exhibited exceptional image-text joint understanding and generalization abilities, facilitating a large spectrum of downstream applications. Among the works major in MLLMs, LLaVA (Liu et al., 2024b; 2023) and MiniGPT4 (Zhu et al., 2023; Chen et al., 2023a) observed that multimodal SFT is sufficient to align MLLMs with human preferences and enable them to accurately answer fine-grained questions about visual content. Besides, Video-LLaMA (Zhang et al., 2023b) and VideoChat (Li et al., 2023) utilized MLLMs for video understanding. VILA (Lin et al., 2023) quantitatively proved that involving text-only instruction-tuning data during SFT can further ameliorate model performance on text-only and multimodal downstream tasks. LLaVA-NeXT (Liu et al., 2024a) extracted visual tokens for both the resized input image and the segmented sub-images to provide more detailed visual information for MLLMs, achieving significant performance bonus on tasks with high-resolution input images.
However, due to the data bias, existing MLLMs cannot perfectly quantify for text-to-image evaluations. Thus, we meticulously curate a SFT dataset to align MLLMs with detailed human feedback on synthesized images.
@@ 3 EvalAlign Dataset Construction
To train, validate and test the effectiveness of our evaluation models, we build EvalAlign dataset. Specifically, EvalAlign dataset is a meticulously annotated collection featuring fine-grained annotations for images generated on text conditions. This dataset comprises 21k images, each accompanied by detailed instructions. The compilation process for the EvalAlign Dataset encompasses prompt collection, image generation, and precise instruction-based annotation.
@@ 3.1 Prompts and Images Collection
Prompt collection. To assess the capabilities of our model in terms of image faithfulness and text-image alignment, we collect, filter, and clean prompts from existing evaluation datasets and generated prompts based on LLM. These prompts encompass a diverse range from real-world user prompts, prompts generated through rule-based templates with LLM, to manually crafted prompts. Specifically, the utilized prompts are soureced from HPS (Wu et al., 2023b), HRS-Bench (Bakr et al., 2023), HPSv2 (Wu et al., 2023a), TIFA (Hu et al., 2023), DSG (Cho et al., 2023a), T2I-Comp (Huang et al., 2023a), Winoground (Thrush et al., 2022), DALL-EVAL (Cho et al., 2023b), DiffusionDB (Wang et al., 2023), PartiPrompts (Yu et al., 2022), DrawBench (Saharia et al., 2022), and JourneryDB (Sun et al., 2024).
Prompt curation. To facilitate a clean and reasonable evaluation, each prompt to be annotated have to instruct text-to-image models to generate images that can reflect model performances on image faithfulness and text-image alignment. However, considering some of the collected prompts fail to achieve the purpose, we need to filter and balance the collected prompts to ensure their quantity, quality and diversity. For image faithfulness evaluation, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment. Consequently, the prompt filter for image faithfulness initially selects prompts that describe human, animals, and other real objects. After deduplicating these prompts, we carefully select 1,500 distinct prompts with varying topic, background and style. The selected prompts encompass 10k subjects across 15 categories. For text-image alignment evaluation, we refine our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts. Specifically, only prompts contain relevant descriptions and exceed 15 words in length are considered, culminating in a final set of 1,500 prompts.
Image generation. To train and evaluate the MLLM, we use a diverse set of images generated by various models using the aforementioned prompts, facilitating detailed human annotation. For each prompt, multiple images are generated across different models. The models used to generate these images vary in architectures and scales, enhancing the dataset diversity. There are 24 models used to generate these images, varying in architecture as well as scale and thus enhancing the dataset diversity. For detailed information on the generation setting of each model, please refer to the appendix.
The training and validation set comprises synthesized images from 8 out of the 24 models, whereas the test set spans all of them. Particularly, the exclusive inclusion of the 16 models in the test set is crucial for validating the MLLM’s ability to generalize beyond its training data. Through our manual inspection, in this way, we attain ample synthesized images with a balanced diversity in the performance of image faithfulness and text-image alignment.
@@ 3.2 Data Annotation
Prompt annotation. For text prompts focused on text-image alignment, we begin by annotating the entities and their attributes within the text, as illustrated in Figure 1. Our annotators extract the entities mentioned in the prompts and label each entity with corresponding attributes, including quantity, color, spatial relationships, and actions. During the annotation, we also ask the annotators to annotate the overall style of the image if described in the corresponding prompt and report prompts that contain toxic and NSFW content. These high-quality and detailed annotations facilitate the subsequent SFT training and evaluation of the MLLM. The prompt annotation procedure ensures that the MLLM can accurately align and respond to the nuanced details specified in the prompts, enhancing both the training process and the model’s performance in generating images that faithfully reflect the described attributes and style.
@@
Image annotation. The images generated by text-to-image models often present challenges such as occluded human body parts, which can impede the effectiveness of SFT training and evaluation of the MLLM. To address these challenges and enhance the model’s training and evaluative capabilities, specific annotations are applied to all images depicting human and animals. These annotations include: presence of human or animal faces; visibility of hands; visibility of limbs. By implementing these annotations, we ensure that the MLLM can more effectively learn from and assess the completeness and faithfulness of the generated images. This structured approach to annotation not only aids in identifying common generation errors but also optimizes the model’s ability to generate more accurate and realistic images, thereby improving both training outcomes and the model’s overall performance in generating coherent and contextually appropriate visual content.
@@
Instruction-finetuning data annotation. To align the MLLM with human preference prior on detailed synthesized image assessing, we can train the model on a minimal amount of fine-grained human feedback data through SFT training. As a consequence, we devise two sets of questions, each is concentrated on a specific fine-grained skill of image faithfulness and image-text alignment. Human annotators are required to answer these questions to acquire the fine-grained human preference data. To aid them to understand the meaning and principle of each question, thereby ensuring high annotation quality, we employ a thorough and comprehensive procedure of annotation preparation. First, we write a detailed annotation guideline and conduct a training for the annotators to explain the annotation guideline and answer their questions about the annotation. Then, we conduct a multi-turn trial annotation on another 50 synthesized images. After each trial, we calculate the Cohen’s kappa coefficient and interpret annotation guidelines for our annotators. In total, we conduct nine turns of trial annotation, and in the last turn of the trial, the Cohen’s kappa coefficient of our annotators reaches , indicating high inter-annotator reliability and high annotation quality.
@@
After completing the aforementioned preparations, we delegate the images filtered during image annotation to 10 annotators and ask them to complete the annotation just as how they did in the trial annotation. Furthermore, during the whole annotation procedure, four experts in text-to-image generation conduct random sampling quality inspection on the present annotated results, causing a second and a third re-annotation on 423 and 112 inspection-failed samples. Overall, owing to the valuable work of our human annotators and our fastidious annotation procedure, we get quality-sufficient instruction-tuning data required for the SFT training of the MLLM. More details of the annotation procedure will be introduced in supplementary files.